Please answer all and show your work

Assume Gasol Company has direct labor costs of $1,096, manufacturing overhead costs of $1,457 and prime costs of $3,242. What is Gasol's direct materials cost?

-------------------------

Assume Gasol Company has direct materials cost of $2,661, direct labor costs of $1,018, manufacturing overhead costs of $1,362. What is Gasol's prime cost?

------------------------

Leonard, Inc., which uses a volume-based cost system, produces cat condos that sell for $104 each. Direct materials cost $11 per unit, and direct labor costs $23 per unit. Manufacturing overhead is applied at a rate of 180% of direct labor cost. Nonmanufacturing costs are $18 per unit. What is the gross profit per unit for the cat condos? (Round your answer nearest whole dollar.)

this is the question as it is posted.

----------------------

Job 101 has a direct materials cost of $1,100 and a total manufacturing cost of $2,600. Overhead is applied to jobs at a rate of 100 percent of direct labor cost.

What is the direct labor cost?

-----------------------
Answer #1

a) Direct material cost = Prime cost-Direct labor = 3242-1096 = 2146

b) Prime cost = Direct material+Direct labor = 2661+1018 = 3679

c) Gross profit per unit = 104-(11+23+23*1.8) = 28.60

d) Direct labor cost =

Direct labor and overhead = 2600-1100 = 1500

Direct labor = 1500/2 = 750